Session: Mar 2010 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (C.Bortolotti at Sat Mar 06 21:42:21 2010) Success:

Noto (P. Cassaro at Sun Mar 07 11:33:22 2010) Minor failures: TSYS
BBC #8 have no tsys

Onsala (Michael Lindqvist at Sun Mar 07 14:02:31 2010) Success:
No known problems.

Effelsberg (dg at Mon Mar 08 10:00:27 2010) Minor failures: RECO LATE
late start 16:03 due to mk5 problem (slow pack)

Torun (Kaz Boerkowski at Mon Mar 08 13:14:07 2010) Success: RECO
Recording started 16 minutes late due to a problem with MK5 (operator had to reboot it).

Shanghai (Bo Xia at Tue Mar 09 07:09:34 2010) Success:

Noto (P. Cassaro at Wed Mar 10 11:27:07 2010) Minor failures:
ADDENDUM: performed with the rubidium clock

Jodrell_Bank (alastair gunn at Mon Mar 29 10:52:24 2010) Minor failures: ANT
Frequent telescope encoder faults giving brief periods off source. No other known problems.

Cambridge (alastair gunn at Mon Mar 29 10:53:28 2010) Success:
Kn on BBCs 1 and 2. No known problems.